A test has 20 questions worth 100 points. The test consists of true and false questions worth 3 points each and multiple choice questions worth 11 points each. How many true and false and multiple choice questions are on
the test

Answer by TimothyLamb(4379)

---
x = t/f
y = m/c
---
x + y = 20
3x + 11y = 100
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = t/f = 15
y = m/c = 5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php